If 4 boys can clean the gymnasium in 2 hours, how many minutes will it take 5 boys to do the job?
Answer by mananth(16946)

4 boys ------------- 2 hours
5 boys -------------x hours
indirect variation
5/4=2/x
x=8/5
..
8/5 * 60 = 96 minutes
...